Perimeter of a rectangle. The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find a binomial that represents the length. please show all work and solution. Thanks
Answer by tutorcecilia(2152)

P=2[length + width] [Formula for a perimeter]
(6x+6)=2[(length) + (x)] [Plug-in the given values]
6x+6=2(length) + 2x [Distribute the 2]
(6x+6) - 2x = 2(length) +2x-2x [Solve for length]
6x+6-2x=2(length)
4x+6=2(length)
(4x+6)/2=2(length)/2
(4x+6)/2=length
2(2x+3)/2 = length [Factor and cancel out]
2x+3=length [Final answer]
.
Check the answer by plugging (length = 2x+3) back into the original formula.
P=2[length + width] [Returning to the formula for a perimeter]
P = 2[(x) + (2x+3)]
P = 2x+4x + 6
P = 6x+6 [Checks out]